# LPO: Towards Accurate GUI Agent Interaction via Location Preference Optimization

### v2

@@ Abstract
The advent of autonomous agents is transforming interactions with Graphical User Interfaces (GUIs) by employing natural language as a powerful intermediary. Despite the predominance of Supervised Fine-Tuning (SFT) methods in current GUI agents for achieving spatial localization, these methods face substantial challenges due to their limited capacity to accurately perceive positional data. Existing strategies, such as reinforcement learning, often fail to assess positional accuracy effectively, thereby restricting their utility. In response, we introduce Location Preference Optimization (LPO), a novel approach that leverages locational data to optimize interaction preferences. LPO uses information entropy to predict interaction positions by focusing on zones rich in information. Besides, it further introduces a dynamic location reward function based on physical distance, reflecting the varying importance of interaction positions. Supported by Group Relative Preference Optimization (GRPO), LPO facilitates an extensive exploration of GUI environments and significantly enhances interaction precision. Comprehensive experiments demonstrate LPO’s superior performance, achieving SOTA results across both offline benchmarks and real-world online evaluations. Our code will be made publicly available soon, at https://github.com/AIDC-AI/LPO.
@@ 1 Introduction
The advent of autonomous agents has profoundly altered strategies for Graphical User Interface (GUI) interactions [26, 11, 21]. By utilizing natural language as an intermediary [8], these agents minimize labor and time costs associated with manual GUI operations, thus leading to their growing prevalence in recent times [26].
@@
Most GUI agents rely heavily on Supervised Fine-Tuning (SFT) during the training process [8, 3, 2, 7]. However, SFT often encounters significant challenges in spatial localization due to its limited capability to perceive and interpret positional data [18]. This shortcoming impairs precise interactions within the GUI, highlighting the fundamental challenge of improving the accuracy of such interactions.
@@
Despite some strategies [18, 23, 16, 28, 13] attempting to utilize Reinforcement Learning (RL) to enhance the accuracy of UI action decisions, these methods often lack a mechanism for accurately assessing interactions’ positional accuracy. As a result, their ability to improve interaction accuracy is limited (as illustrated in Figure 1 (a) & (b) & (c)). Additionally, some methods like UI-TARS [18] rely heavily on manually constructing positive and negative actions for direct preference optimization, thereby becoming highly dependent on data construction. Consequently, these methods fail to fully resolve the issue of precise spatial localization during GUI interactions.
@@
To align precise GUI interaction, we introduce Location Preference Optimization (LPO), an innovative approach that leverages locational data for optimizing accurate interaction preferences. Specifically, drawing inspiration from the tendency of users to interact more frequently in zones with higher information density, we divide the interface into distinct windows and employ their information entropy to build a reward for preliminarily forecasting interaction positions (see Section 4.1). Subsequently, to offer a more nuanced representation of the varying significance of interaction positions, we incorporate physical distance to develop a dynamic location reward function (see Section 4.2 and Figure 1 (d)). Finally, by integrating these rewards, we implement LPO, inspired by Group Relative Preference Optimization (GRPO) [19]. This methodology enables a more comprehensive exploration of expansive GUI environments, guiding the agent to optimize preferences that correspond to precise interaction capabilities (see Section 4.3).
@@
Our experimental results comprehensively demonstrate that LPO significantly enhances the interaction capabilities of GUI agents, achieving state-of-the-art (SOTA) performance compared to other preference optimization strategies. This improvement is evident in offline benchmarks, both in GUI Interaction (Multimodal Mind2Web [4]) and Grounding (VisualWebBench [12] and Screenspot V2 [22]). Furthermore, our approach also exhibits superior performance in real-world scenarios during online evaluations (WebVoyager [7]).
@@
Figure 1: Motivation of dynamic location reward. (a) UITARS [18] uses direct text-level matching; (b) UI-R1 [16], InfiGUI-R1 [13] and RUIG [28] employ bounding boxes for interaction preferences; (c) GUI-R1 [23] relies on fixed positional boundaries. (d) Our dynamic location reward offers a more precise positional representation, addressing the limitations of previous methods.
@@ 2 Related Work
The development of Multimodal Large Language Models (MLLMs) has recently empowered users to create GUI Agents capable of automating interactions with user interfaces to meet specific user demands  [15, 18, 8]. Nevertheless, determining the optimal strategy for facilitating accurate interaction between agents and GUIs remains a significant challenge.
Early approaches utilizing Set-of-Mark (SoM) identified candidate buttons and click locations on graphical interfaces [24]. Despite their functionality, these methods limited decision space and were prone to missed or false detections, causing interaction inaccuracies. Besides, some solutions attempted to interact directly through raw source code (e.g., HTML, APIs) [5, 17], but these approaches lack intuitive visual grounding, hindering natural graphical interface interaction. Most recently, the interaction mode has shifted focus to vision-based strategies, allowing agents to use visual inputs and text outputs for GUI operations [8]. This approach bypasses earlier constraints by letting agents analyze interface regions freely and align with visual elements intuitively.
@@
The accurate grounding ability of GUI agents, based on visual perception, is crucial for precise interaction. Recently, methods such as those by Gou et al. [6] and Cheng et al. [2] have attempted to learn GUI grounding capabilities directly through Supervised Fine-Tuning (SFT). However, this process often involves challenges related to data format alignment and unclear physical information, making it difficult to achieve more precise localization performance.
@@
Recently, various preference optimization strategies have emerged as significant tools in GUI Agents. Qin et al. [18] introduced Direct Preference Optimization (DPO) using positive and negative samples from interaction paths to amend erroneous interactions. However, this requires manual construction of sample pairs, which can be labor-intensive and limiting. Xia et al. [23] and Lu et al. [16] developed Rule-based Preference Optimization to assess the accuracy of predicted interaction actions. In contrast, Zhang et al. [28] and Liu et al. [13] employed bounding box positions with fixed threshold constraints to differentiate positive and negative examples. Despite their effectiveness in evaluating interaction accuracy, these approaches commonly rely on static decision boundaries, which offer only coarse evaluations of spatial relationships, leading to imprecise interaction localization.
@@ 4 Methodology
To achieve more precise GUI interactions, although previous approaches [16, 23, 13] have utilized physical rewards based on interaction space (e.g., IoU or fixed decision boundary), the assessment of rewards for positions remains imprecise (as discussed in Section 2).
@@
In this paper, we propose Location Preference Optimization (LPO), a novel approach that precisely leverages accurate locational data for preference optimization. Firstly, considering that users are more inclined to interact in zones with higher information densities, we segment the interface into distinct windows and utilize their information entropy for a preliminary forecast of interaction positions (Section 4.1). Secondly, to provide a finer representation of varying importance across interaction positions, we utilize physical distance to construct a location-based reward metric (Section 4.2). Lastly, by amalgamating these rewards, we introduce LPO, grounded in the Group Relative Preference Optimization (GRPO) [19]. This approach facilitates a more broader exploration of expansive GUI spaces and directs the agent to optimize towards preferences aligned with precise interaction capabilities (Section 4.3).
Figure 2: Example of $r_{w}$. Green zones indicate high interaction likelihood due to rich information, earning greater rewards. In contrast, red zones, like blank areas, have lower interaction probability and rewards. Key interactive areas, such as login, search, and editing zones, align with user interaction tendencies.
Figure 3: Example of $r_{d}$. When users need to interact at a point located on the search button, the reward increases as the generated interaction point gets closer to this target point, while it decreases as the point moves further away. This highlights the importance of precision in interaction positioning.
@@ 4.1 Window-based Information Density Reward
In GUI interaction tasks, an agent iteratively observes the current visual state $s_{t}\in\mathbb{R}^{C\times H\times W}$, executes an action $a_{t}\in(\mathcal{A}_{t}\times\mathcal{E}_{t})$, and transitions to the subsequent state $s_{t+1}$ following the trajectory $s_{t}\rightarrow a_{t}\rightarrow s_{t+1}$. Visual states exhibit heterogeneous information density across spatial regions, where functional elements (e.g., buttons, text fields) predominantly occupy high-density zones. To enhance the precision of the agent’s interactions within these critical regions, we propose a reward mechanism designed to incentivize actions targeting these zones.
@@
$\mathbf{W}_{i,j}=s_{t}\left[:,\frac{(i-1)H}{M}:\frac{iH}{M},\frac{(j-1)W}{N}:\frac{jW}{N}\right],\quad\forall i\in\{1,\ldots,M\},\ j\in\{1,\ldots,N\},$ (3)
where $\mathbf{W}_{i,j}$ denotes the window at grid position $(i,j)$. To ensure consistent visual perceptual capacity across the windows in one image, we empirically maintain $M$ and $N$ to match the same settings used by multi-modal large language models.
@@
where $B$ is the total number of bins in the histogram. This entropy measure quantifies the amount of information or uncertainty present in the window’s pixel intensity distribution across all grid positions $(i,j)$.
@@
$r_{w}=\frac{\mathcal{H}_{i^{*},j^{*}}}{\max\limits_{i,j}\mathcal{H}_{i,j}+\epsilon},\quad\text{where }\begin{cases}i^{*}=\lceil\frac{y}{H/M}\rceil\\ j^{*}=\lceil\frac{x}{W/N}\rceil\end{cases},$ (5)
@@ 4.2 Dynamic Location Reward
[…]
where $d_{max}$ represents the maximum allowable distance used for scaling the reward, set at $1000$.
@@
$r_{d}=\begin{cases}\frac{1}{K}\sum_{k=1}^{K}r_{k},&\text{if }\mathcal{A}_{t}=\mathcal{A}^{*}\\ 0,&\text{otherwise}\end{cases},$ (7)
@@ 4.3 Location Preference Optimization
To explore a broader space in GUI, based on GRPO [19], we leverage our location-based reward functions to measure relative location advantages. Our advantage definition is formulated as in Eq. 8,
$A^{(g)}=\frac{r^{(g)}-\text{mean}(\sum_{g=1}^{G}r^{(g)})}{\text{std}(\sum_{g=1}^{G}r^{(g)})},\quad r^{(g)}=r_{w}^{(g)}\times r_{d}^{(g)},$ (8)
where $r_{w}^{(g)}$ and $r_{d}^{(g)}$ represent the rewards in $g$-times exploitation, and $G$ is the group size. $A^{(g)}$ is the advantage that emphasizes relative position comparison.
@@
$\small\begin{split}\mathcal{J}_{\text{LPO}}&(\theta)=\mathbb{E}_{\begin{subarray}{c}\{a_{g}\}_{g=1}^{G}\sim\pi_{\theta_{\text{old}}}\end{subarray}}\\ &\frac{1}{G}\sum_{v=1}^{G}\Bigg{[}\min\bigg{(}\underbrace{\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}}_{\text{Importance Ratio}}A^{(g)},\text{clip}\Big{(}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})},1-\epsilon_{1},1+\epsilon_{2}\Big{)}A^{(g)}\bigg{)}-\beta\underbrace{\mathbb{D}_{\text{KL}}\big{(}\pi_{\theta}\parallel\pi_{\text{ref}}\big{)}}_{\text{KL Regularization}}\Bigg{]},\end{split}$ (9)

### v3

@@ Abstract
The advent of autonomous agents is transforming interactions with Graphical User Interfaces (GUIs) by employing natural language as a powerful intermediary. Despite the predominance of Supervised Fine-Tuning (SFT) methods in current GUI agents for achieving spatial localization, these methods face substantial challenges due to their limited capacity to accurately perceive positional data. Existing strategies, such as reinforcement learning, often fail to assess positional accuracy effectively, thereby restricting their utility. In response, we introduce Location Preference Optimization (LPO), a novel approach that leverages locational data to optimize interaction preferences. LPO uses information entropy to predict interaction positions by focusing on zones rich in information. Besides, it further introduces a dynamic location reward function based on physical distance, reflecting the varying importance of interaction positions. Supported by Group Relative Preference Optimization (GRPO), LPO facilitates an extensive exploration of GUI environments and significantly enhances interaction precision. Comprehensive experiments demonstrate LPO’s superior performance, achieving SOTA results across both offline benchmarks and real-world online evaluations. Our code will be made publicly available soon, at https://github.com/jqtangust/LPO.
LPO: Towards Accurate GUI Agent Interaction via
Location Preference Optimization
Jiaqi Tang1111Equal contribution   Yu Xia2111Equal contribution   Yi-Feng Wu2111Equal contribution   Yuwei Hu2111Equal contribution   Yuhui Chen2   Qing-Guo Chen2 Xiaogang Xu3222Corresponding authors: Qifeng Chen (cqf@ust.hk) and Xiaogang Xu (xiaogangxu00@gmail.com).Xiangyu Wu4Hao Lu5Yanqing Ma2Shiyin Lu2Qifeng Chen1222Corresponding authors: Qifeng Chen (cqf@ust.hk) and Xiaogang Xu (xiaogangxu00@gmail.com). 1The Hong Kong University of Science and Technology  2Alibaba Group 3The Chinese University of Hong Kong  4Nanjing University of Science and Technology 5The Hong Kong University of Science and Technology (Guangzhou)
@@ 1 Introduction
The advent of autonomous agents has profoundly altered strategies for Graphical User Interface (GUI) interactions Zhang et al. (2024a); Lieberman (1997); Wang et al. (2024). By utilizing natural language as an intermediary Hong et al. (2023), these agents minimize labor and time costs associated with manual GUI operations, thus leading to their growing prevalence in recent times Zhang et al. (2024a).
@@
Most GUI agents rely heavily on Supervised Fine-Tuning (SFT) during the training process Hong et al. (2023); Deng et al. (2023a); Cheng et al. (2024); He et al. (2024). However, SFT often encounters significant challenges in spatial localization due to its limited capability to perceive and interpret positional data Qin et al. (2025). This shortcoming impairs precise interactions within the GUI, highlighting the fundamental challenge of improving the accuracy of such interactions.
Figure 1: Motivation of dynamic location reward. (a) UITARS Qin et al. (2025) uses direct text-level matching; (b) UI-R1 Lu et al. (2025), InfiGUI-R1 Liu et al. (2025) and RUIG Zhang et al. (2023) employ bounding boxes for interaction preferences; (c) GUI-R1 Xia and Luo (2025) relies on fixed positional boundaries. (d) Our dynamic location reward offers a more precise positional representation, addressing the limitations of previous methods.
@@
Despite some strategies Qin et al. (2025); Xia and Luo (2025); Lu et al. (2025); Zhang et al. (2023); Liu et al. (2025) attempting to utilize Reinforcement Learning (RL) to enhance the accuracy of UI action decisions, these methods often lack a mechanism for accurately assessing interactions’ positional accuracy. As a result, their ability to improve interaction accuracy is limited (as illustrated in Figure 1 (a) & (b) & (c)). Additionally, some methods like UI-TARS Qin et al. (2025) rely heavily on manually constructing positive and negative actions for direct preference optimization, thereby becoming highly dependent on data construction. Consequently, these methods fail to fully resolve the issue of precise spatial localization during GUI interactions.
@@
To align precise GUI interaction, we introduce Location Preference Optimization (LPO), an innovative approach that leverages locational data for optimizing accurate interaction preferences. Specifically, drawing inspiration from the tendency of users to interact more frequently in zones with higher information density, we divide the interface into distinct windows and employ their information entropy to build a reward for preliminarily forecasting interaction positions (see Section 4.1). Subsequently, to offer a more nuanced representation of the varying significance of interaction positions, we incorporate physical distance to develop a dynamic location reward function (see Section 4.2 and Figure 1 (d)). Finally, by integrating these rewards, we implement LPO, inspired by Group Relative Preference Optimization (GRPO) Shao et al. (2024). This methodology enables a more comprehensive exploration of expansive GUI environments, guiding the agent to optimize preferences that correspond to precise interaction capabilities (see Section 4.3).
@@
Our experimental results comprehensively demonstrate that LPO significantly enhances the interaction capabilities of GUI agents, achieving state-of-the-art (SOTA) performance compared to other preference optimization strategies. This improvement is evident in offline benchmarks, both in GUI Interaction (Multimodal Mind2Web Deng et al. (2023b)) and Grounding (VisualWebBench Liu et al. (2024) and Screenspot V2 Wu et al. (2024)). Furthermore, our approach also exhibits superior performance in real-world scenarios during online evaluations (WebVoyager He et al. (2024)).
@@ 2 Related Work
The development of Multimodal Large Language Models (MLLMs) Tang et al. (2024a, 2026b, 2026a); Lu et al. (2024a) has recently empowered users to create GUI Agents capable of automating interactions with user interfaces to meet specific user demands  Lu et al. (2024c); Qin et al. (2025); Hong et al. (2023). Nevertheless, determining the optimal strategy for facilitating accurate interaction between agents and GUIs remains a significant challenge.
Early approaches utilizing Set-of-Mark (SoM) identified candidate buttons and click locations on graphical interfaces Yang et al. (2023). Despite their functionality, these methods limited decision space and were prone to missed or false detections, causing interaction inaccuracies. Besides, some solutions attempted to interact directly through raw source code (e.g., HTML, APIs) Furuta et al. (2024); Lù et al. (2024), but these approaches lack intuitive visual grounding, hindering natural graphical interface interaction. Most recently, the interaction mode has shifted focus to vision-based strategies, allowing agents to use visual inputs and text outputs for GUI operations Hong et al. (2023). This approach bypasses earlier constraints by letting agents analyze interface regions freely and align with visual elements intuitively.
@@
The accurate grounding ability of GUI agents, based on visual perception Tang et al. (2024a, b, c), is crucial for precise interaction. Recently, methods such as those by Gou et al. Gou et al. (2025) and Cheng et al. Cheng et al. (2024) have attempted to learn GUI grounding capabilities directly through Supervised Fine-Tuning (SFT). However, this process often involves challenges related to data format alignment and unclear physical information, making it difficult to achieve more precise localization performance.
@@
Recently, various preference optimization strategies have emerged as significant tools in GUI Agents. Qin et al. Qin et al. (2025) introduced Direct Preference Optimization (DPO) using positive and negative samples from interaction paths to amend erroneous interactions. However, this requires manual construction of sample pairs, which can be labor-intensive and limiting. Xia et al. Xia and Luo (2025) and Lu et al. Lu et al. (2025) developed Rule-based Preference Optimization to assess the accuracy of predicted interaction actions. In contrast, Zhang et al. Zhang et al. (2023) and Liu et al. Liu et al. (2025) employed bounding box positions with fixed threshold constraints to differentiate positive and negative examples. Despite their effectiveness in evaluating interaction accuracy, these approaches commonly rely on static decision boundaries, which offer only coarse evaluations of spatial relationships, leading to imprecise interaction localization.
@@ 4 Methodology
To achieve more precise GUI interactions, although previous approaches Lu et al. (2025); Xia and Luo (2025); Liu et al. (2025) have utilized physical rewards based on interaction space (e.g., IoU or fixed decision boundary), the assessment of rewards for positions remains imprecise (as discussed in Section 2).
@@
In this paper, we propose Location Preference Optimization (LPO), a novel approach that precisely leverages accurate locational data for preference optimization. Firstly, considering that users are more inclined to interact in zones with higher information densities, we segment the interface into distinct windows and utilize their information entropy for a preliminary forecast of interaction positions (Section 4.1). Secondly, to provide a finer representation of varying importance across interaction positions, we utilize physical distance to construct a location-based reward metric (Section 4.2). Lastly, by amalgamating these rewards, we introduce LPO, grounded in the Group Relative Preference Optimization (GRPO) Shao et al. (2024). This approach facilitates a more broader exploration of expansive GUI spaces and directs the agent to optimize towards preferences aligned with precise interaction capabilities (Section 4.3).
Figure 2: Example of $r_{\mathrm{w}}$. Green zones indicate high interaction likelihood due to rich information, earning greater rewards. In contrast, red zones, like blank areas, have lower interaction probability and rewards. Key interactive areas, such as login, search, and editing zones, align with user interaction tendencies.
@@ 4.1 Window-based Information Density Reward
In GUI interaction tasks, an agent iteratively observes the current visual state $s_{t}\in\mathbb{R}^{C\times H\times W}$, executes an action $a_{t}\in(\mathcal{A}_{t}\times\mathcal{E}_{t})$, and transitions to the subsequent state $s_{t+1}$ following the trajectory $s_{t}\rightarrow a_{t}\rightarrow s_{t+1}$. The distribution of visual information across the interface is heterogeneous. Functional elements like buttons and text fields cluster in regions of high information density. To steer the agent’s focus towards these critical regions, we introduce a window-based information density reward.
@@
$\mathbf{W}_{i,j}$ $=s_{t}\left[:,\frac{(i-1)H}{M}:\frac{iH}{M},\frac{(j-1)W}{N}:\frac{jW}{N}\right],$ (3)
$\quad\forall i\in\{1,\ldots,M\},\ j\in\{1,\ldots,N\},$
where $\mathbf{W}_{i,j}$ denotes the window at grid position $(i,j)$. To ensure consistent visual perceptual capacity across the windows in one image, we empirically set $M$ and $N$ to align with the visual tokenization pattern of the underlying based multimodal large language model Lu et al. (2024b), ensuring that each window corresponds to the same spatial resolution as the model’s tokenizer segmentation scheme.
@@
where $\mathbf{W}_{i,j}$ denotes the window at grid position $(i,j)$. To ensure consistent visual perceptual capacity across the windows in one image, we set $M$ and $N$ to match the visual tokenizer’s patch segmentation scheme of the underlying multi-modal large language model Lu et al. (2024b). Specifically, we align our window dimensions such that each window corresponds to the same spatial resolution as the visual patches, i.e., $M=\lceil H/h\rceil$ and $N=\lceil W/w\rceil$, where $H$ and $W$ are the predefined height and width, respectively.
@@
$r_{\mathrm{w}}=\frac{\mathcal{H}_{i^{*},j^{*}}}{\max\limits_{i,j}\mathcal{H}_{i,j}+\epsilon},\ \text{where }\begin{cases}i^{*}=\lceil\frac{y}{H/M}\rceil\\ j^{*}=\lceil\frac{x}{W/N}\rceil\end{cases},$ (5)
@@ 4.2 Dynamic Location Reward
While the window-based reward encourages exploration of information-rich regions, precise task execution also requires accurate targeting of specific coordinates. To this end, we introduce a dynamic location reward that directly measures spatial accuracy.
@@
Figure 3: Example of $r_{\mathrm{d}}$. When users need to interact at a point located on the search button, the reward increases as the generated interaction point gets closer to this target point, while it decreases as the point moves further away. This highlights the importance of precision in interaction positioning.
@@
$r_{k}$ $=\max\left(0,\ 1-\frac{\sqrt{(x^{k}-x^{*k})^{2}+(y^{k}-y^{*k})^{2}}}{d_{\max}}\right),$ (6)
$\quad\forall k\in\{1,\dots,K\},$
where $d_{\max}$ denotes the maximum distance used to normalize the reward; we set $d_{\max}=1000$.
@@
$r_{\mathrm{d}}=\begin{cases}\frac{1}{K}\sum_{k=1}^{K}r_{k},&\text{if }\mathcal{A}_{t}=\mathcal{A}^{*}\\ 0,&\text{otherwise}\end{cases},$ (7)
@@ 4.3 Location Preference Optimization
To explore a broader space in GUI, based on GRPO Shao et al. (2024), we leverage our location-based reward functions to measure relative location advantages. Our advantage definition is formulated as in Eq. 8,
$\begin{gathered}A^{(g)}=\frac{r^{(g)}-\text{mean}(\sum_{g=1}^{G}r^{(g)})}{\text{std}(\sum_{g=1}^{G}r^{(g)})},\\ r^{(g)}=r_{\mathrm{w}}^{(g)}\,r_{\mathrm{d}}^{(g)},\end{gathered}$ (8)
where $r_{\mathrm{w}}^{(g)}$ and $r_{\mathrm{d}}^{(g)}$ denote the two reward factors for the $g$-th sample in each group, and $G$ is the group size. $A^{(g)}$ is the advantage that emphasizes relative position comparison.
@@
$\mathcal{J}_{\text{LPO}}(\theta)$ $=\mathbb{E}_{\begin{subarray}{c}\{a_{g}\}_{g=1}^{G}\sim\pi_{\theta_{\text{old}}}\end{subarray}}$ (9)
$\frac{1}{G}\sum_{v=1}^{G}\Bigg[$ $\quad\min\bigg(\underbrace{\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}}_{\text{Importance Ratio}}A^{(g)},$
$\quad\text{clip}\Big(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})},1-\epsilon_{1},1+\epsilon_{2}\Big)A^{(g)}\bigg)$
$\quad-\beta\underbrace{\mathbb{D}_{\text{KL}}\big(\pi_{\theta}\parallel\pi_{\text{ref}}\big)}_{\text{KL Regularization}}\Bigg],$
@@
Table 1: Performance of GUI interaction on Multimodal Mind2Web Deng et al. (2023b). We report Element Accuracy (Ele.Acc), Operation F1 (Op.F1) and Step Success Rate (Step SR). The best model is in-bold, and the second best is underlined.
Ele.Acc Op.F1 Step SR Ele.Acc Op.F1 Step SR Ele.Acc Op.F1 Step SR
After Supervised Fine-Tuning
Base Model $60.3$ $57.4$ $38.2$ $60.7$ $56.9$ $38.4$ $63.8$ $58.5$ $40.7$
After Preference Optimization
+ $R_{\text{UI-R1}}$ Lu et al. (2025) $59.5$ $34.5$ $24.9$ $56.5$ $31.5$ $22.1$ $61.6$ $37.2$ $27.1$
+ $R_{\text{GUI-R1}}$ Xia and Luo (2025) $62.5$ $\underline{71.6}$ $\underline{46.6}$ $61.2$ $67.6$ $43.5$ $65.0$ $71.1$ $47.9$
+ $R_{\text{InfiGUI-R1}}$ Liu et al. (2025) $\underline{62.6}$ $51.3$ $35.8$ $\underline{62.2}$ $49.5$ $34.4$ $\underline{65.1}$ $53.1$ $40.0$